# Tensor network methods for quantum-inspired image processing and classical optics

### v1

@@ Abstract
Tensor network methods strike a middle ground between fully-fledged quantum computing and classical computing, as they take inspiration from quantum systems to significantly speed up certain classical operations. Their strength lies in their compressive power and the wide variety of efficient algorithms that operate within this compressed space. In this work, we focus on applying these methods to fundamental problems in image processing and classical optics such as wave-front propagation and optical image formation, by using directly or indirectly parallels with quantum mechanics and computation. These quantum-inspired methods are expected to yield faster algorithms with applications ranging from astronomy and earth observation to microscopy and classical imaging more broadly.
@@ I Introduction
Recently, the field of tensor networks has seen a surge in interest due to their connections to quantum computing. Quantum researchers use tensor networks to design and simulate their quantum machines, while tensor networks experts borrow ideas from quantum theory to emulate and compete with quantum performance Huang et al. (2019); Oh et al. (2023); Liu et al. (2023); Tindall et al. ; Patra et al. (2024). This feedback loop benefits both sides and benefits also scientists and engineers tackling complex real-world problems. To name a few examples, tensor networks have enabled immediate applications in machine learning Stoudenmire and Schwab (2016); Glasser et al. (2018), fluid mechanics Peddinti et al. (2024); Arenstein et al. (2025); van Hülst et al. (2025); Gourianov et al. (2025); Hölscher et al. (2025), finance Rihito et al. (2025), electromagnetism Lively et al. (2025), material science Hauck et al. (2024) and computational chemistry Jolly et al. (2025).
@@
Outline of the paper In the first half of this paper, we focus on natural images and quantum-inspired methods for encoding them into quantum-like objects, which will prove useful for many applications. Several types of matrix product encodings will be compared with one another, along with alternative tensor network structures such as tree tensor networks. The discussion will then shift to classical operations on images represented as tensor network operators, and how these can be applied to classical imaging and image processing. One of the goals of this paper is to draw parallels between Fourier optics and quantum mechanics, and to use these parallels to construct efficient approximations of classical optical propagation in very large dimensions by finding local Hamiltonians whose evolution represents these classical operations. These connections will be often made using ideas stemming from quantum computation—even though no quantum computers are involved—showing how quantum mechanics concepts can lead to efficient classical computations.
II Natural images and quantum encodings
We define a continuous ”image” (sometimes called a scene or an objet) as the intensity $h$ of an electromagnetic field of a classical object on a $2d$ plane. From this function of continuous variables, we can consider an approximation on a $2d$ grid of size $L\times L=2^{n}\times 2^{n}$ by evaluating it at discrete points $(x,y)=\Big(\frac{a}{2^{n}},\frac{b}{2^{n}}\Big)$ with $a,b\in\{0,1,2,...,2^{n}-1\}$.
@@ II Natural images and quantum encodings
As we can not get access to a real continuous image of the physical world (or a model of it), we will have to settle for a digital image as a proxy to the real world and try to separate what comes from the physical scene to what comes from the measurement device (the digital camera here). The question we are interested in here, is how much information and correlation is relevant in an image at the continuous limit and how this can used to compress efficiently the image as well as implement faster optical and imaging simulations on classical computers.
@@
Natural images are closer to real physical systems that they are from random samples of a generic distribution, moreover they are known to be scale invariant with a power spectrum that obeys a power-law distribution Ruderman (1994, 1997). In a more recent study Jobst et al. (2024), the authors showed that, over a large dataset of various natural images of various shapes and forms, the behavior of the Fourier amplitudes follows $\langle\hat{h}(k_{x},k_{y})\rangle\sim\mathcal{O}(k_{x}^{-1.1})\mathcal{O}(k_{y}^{-1.07})$. In this present work, we choose to work with a single ”statistically average” image  [^1] As a remark, although the use of ”Lena” as the go-to image in the field of image processing has, understandably so, fallen out of favor in recent years for many reasons, its non-power-law behavior could be added to the list of grievances. (cf. Fig. 1) that is very representative of the power-law behavior of natural images. The picture showed numerous times in this paper follows this behavior honorably well (it is actually closer to $\hat{h}(k)\sim k^{-1}$) and saves us from the time/resource consuming task of averaging over a large dataset of images, with all the generalization and statistical limitations that it entails.
@@
Next, we are mapping this matrix to a higher dimensional objet (a tensor) and then operate compression within this space. The first tensor image encoding that we are considering comes from the field of quantum machine learning, where researchers are interested in efficient ways to feed data to quantum computers. The so-called Flexible Representation of Quantum Images (FRQI) Le et al. (2011) is a quantum encoding scheme that non-trivially represents images by a multi-qubit quantum state $|h\rangle$. More precisely, an image of size $2^{n}\times 2^{n}$ is encoded using a normalized superposition of $2n+1$ qubits
$|h\rangle=\frac{1}{2^{n}}\sum_{i=0}^{2^{2n}-1}|c(x_{i})\rangle\otimes|i\rangle,$ (1)
@@
with $|c(x_{i})\rangle=\cos\big(\frac{\pi}{2}x_{i}\big)|0\rangle+\sin\big(\frac{\pi}{2}x_{i}\big)|1\rangle$, where $x_{i}$ encodes the intensity of the pixel at position $i$ (see Fig. 2), and the basis state $|i\rangle$ represents the pixel’s position in the image. Usually, a row by row flattening of the matrix is used, but more complex paths (see Cataldi et al. (2021) for an application of Hilbert curve to quantum systems) such as Hilbert, Peano or Moore curves can be used  [^2] In Fig. 2, for illustration purpose, a self-avoiding walk is generated by a pivot algorithm Oberdorf et al. (2006).
@@
Figure 2: Matrix to vector reshaping. Below the image, the corresponding vector noted $x_{i}$ with $i=0...2^{2n-1}$.
@@
Intuitively, the path should be as compact as possible such that local correlations in the image remain local upon transformation. In the following, we will use mostly Hilbert curve encoding or will specified otherwise. This encoding allows for efficient quantum image processing by parallel manipulation of all pixel values through unitary operations on the quantum state. This way, we only need $2n+1$ qubits to store the information of $2^{2n}$ grayscale pixel values, which is an exponential reduction to the classical case. Although this encoding requires one more qubit that other encodings, it has been shown to work the best on natural images Jobst et al. (2024).
@@
Thanks to the multi-scale nature of typical images, an other interesting and relevant way of encoding images into tensor is the so-called quantics representation Oseledets (2009); Khoromskij (2011) (first appeared in the quantum computing community Wiesner (1996); Zalka (1998); Latorre (2005)) for the tensors as it naturally encodes scale separation. Each component can be thus written as a rank-2 tensor.
@@
Figure 3: Quantics representation of the linear grid $L=8$. Let us denote $h_{ijk}$ the resulting tensor. In this example, we have $x_{1}\rightarrow h_{000}$, $x_{2}\rightarrow h_{100}$$x_{7}\rightarrow h_{111}$ and $x_{8}\rightarrow h_{001}$.
@@
In $2d$, this means splitting the indices corresponding to spatial coordinates $x$ and $y$ in their binary components $(x^{1}x^{2}x^{3}...x^{n})_{2}$ and $(y^{1}y^{2}y^{3}...y^{n})_{2}$ and rearranging them to form new tensor indices $\mu_{p}=(x^{p}y^{p})_{2}$. This yields a $2n$ order tensor $h^{\mu_{1}\mu_{2}...\mu_{n}}$ where each $\mu_{p}$ corresponds to a particular length scale.  [^3] As an example the value of the image $h$ at coordinate $(x,y)=(43,78)$ on a grid of size $L=128$ is given by the tensor component $h^{00101011001110}$. Technically a full tensor of the form $h^{\mu_{1}\nu_{1}\mu_{2}\nu_{2}...\mu_{n}\nu_{n}}$ contains $2^{n}\times 2^{n}$ parameters, as many as the initial matrix we are considering, only in a different tensor shape. A non-trivial decomposition of this tensor into an interconnected network might reduce the number of parameters from an exponentially to (in the best cases) a logarithmically small number. This generic approach applied to our original image is illustrated in Fig. 4.
@@
Figure 4: Tensor network compression of an image. Here, a generic tensor network is showed for the sake of generality. The approximate symbol will quantify the precision of the compression.
@@
This tensor network decomposition could be used as a way of compressing images for transmission and storage or, as we will see later, as a crucial step in a pipeline of transformations allowing to outperform standard optical simulations or image processing. Many different topologies of tensor networks Bridgeman and Chubb (2017); Biamonte and Bergholm (2017); Orús (2019) have been proposed over the years and became very powerful theoretical and numerical tools that allow to capture complex phenomena regularly found in the fields of condensed matter, high-energy physics or quantum computation Montangero et al. (2018); Collura et al. (2024). In the following section, we focus on the two most common topologies of tensor networks, namely matrix product states and tensor trees and evaluate their compressive capabilities.
III Matrix Product States and Tensor Tree Networks
@@
The specific tensor network we are focusing on first, is called a matrix product state (MPS) in the physics literature or a tensor train (TT) in the mathematical community. The matrix decomposition outlined below (see Fig. 5) is a generalization of the singular value decomposition and can be used to decompose a generic tensor into constituent tensors of lower order.
@@
Figure 5: Illustration of the decomposition of high-order tensor into a matrix product state. The arrow represents a sequential singular value decomposition.
@@
Once the image is encoded into a tensor network, the numerical evaluation of the compressive power versus the reconstruction error is measured and the total number of parameters in the network is called the total dimension $D$. Now, we apply the decomposition to the picture showed earlier and analyze the error of the decomposition with respect to the original image for increasing allowed rank and increasing size. For evaluation sake, we downsize the original image to increasing sizes with an interpolating algorithm to simulate an image with increasingly higher resolutions. The goal is to mirror what we would have done with a typical stat-mech system, by studying the rank of the MPS for various sizes and analyze the large-scale thermodynamic limit.
Let us define the mean squared error between the original image and the decompressed image as
@@
Figure 6: Scaling of the mean squared error of the compression depending on system size for Hilbert curve MPS and quantics tensor trains ($q\mathrm{TT}$).
Additionally, we introduce the inverse compression ratio as $\xi(L)=DL^{-2}$, which is the relevant order parameter for this approximation. Numerical simulations (see Fig. 6) on qubit encoding matrix product states and quantics tensor trains suggest the following finite-size scaling
@@
with $\phi(\xi)\sim\xi^{-\alpha}$ with $\alpha\approx 0.75$ and $\kappa\approx 1.5$ for $\xi\ll 1$. In this efficient regime $\epsilon(L)\sim\mathcal{O}(1)$, meaning that the error is (modulo multiplicative logarithmic corrections) independent of system size (consistent with Jobst et al. (2024)), and results are fairly similar between the two encodings. Conversely, the compression ratio at fixed error increases polynomially $\epsilon=C^{0.75}\times L^{1.5}$, so $C(L)\sim L^{2}$ [^4] This is in sharp contrast with local compressors like JPG, since they possess a maximum compression ratio which is given by the division of the original image in small finite-size squares.
@@
In the overfitting regime, the error goes to zero exponentially quickly at a value $\xi\rightarrow\xi_{\infty}$. We notice that $\xi_{\infty}^{q\mathrm{TT}}<\xi_{\infty}^{\mathrm{MPS}}$ by a big margin (because there is one less qubit needed in the quantics representation than FRQI). This is due to the fact that, by construction of the singular value decomposition, the dimension of a random MPS of $n$ sites is $D_{\mathrm{max}}(n)=8/3\times 2^{n}$ so $\xi_{\infty}^{q\mathrm{TT}}=8/3$ and $\xi_{\infty}^{\mathrm{MPS}}=16/3$ and at very high resolution the error is mostly dominated by pixel noise in the image  [^5] A more fondamental question is whether or not a real image of the physical world (not a digital picture) would be admit a true area-law at any reasonable scale.. Interestingly, the value of $\xi_{\infty}$ is also of certain relevance in the intermediate regime (blue region in Fig. 6) near the compressive threshold $\xi=1$ since a smaller value $\xi_{\infty}$ seems to force the curve to accelerate faster towards this line. This fact allows the $q\mathrm{TT}$ to improve over the Hilbert MPS in this intermediate region. This argument will be made more clear in the following section.
@@
Another important decomposition is the tree tensor network (TTN) where the constituent tensors of the network lie on the nodes of a connected acyclic tree graph. The appeal of tensor tree network comes down to their relative ease of constructions and contractions as well as their greater expressive power and has been recently explored in the context of multivariate function approximations Tindall et al. (2024). Two typical examples of trees are pictured below (see Fig. 7).
@@
Figure 7: Left Binary tensor tree network with 14 physical sites. Right Binary tensor network tree with 8 boundary sites.
@@
In Fig. 8, we plot the scaling of $\xi(L)$ for boundary tensor trees and observe a similar scaling as before. In quantum many-body physics, it is common knowledge that matrix product states can efficiently represents $1d$ quantum systems with area-law entanglement, meaning it can capture only short-range correlations. In contrast, the hierarchical structure of tensor trees can also describe more complex types of entanglement, including some long-range quantum correlations, in higher-dimensional or critical systems where entanglement scaling can exceed that of an area law. In the scenario that we are interested in, the image considered follows an area law (since the bond dimensions of the MPS remain constant up to quite a large precision), and tensor trees only marginally improve the compression ratio but not the scaling of the function $\xi(L)$.

### v2

@@ Abstract
Tensor network methods strike a middle ground between fully-fledged quantum computing and classical computing, as they take inspiration from quantum systems to significantly speed up certain classical operations. Their strength lies in their compressive power and the wide variety of efficient algorithms that operate within this compressed space. In this work, we focus on applying these methods to fundamental problems in image compression and processing and classical optics such as wave-front propagation and optical image formation, by using directly or indirectly parallels with quantum mechanics and computation. These quantum-inspired methods are expected to yield faster algorithms with applications ranging from astronomy and earth observation to microscopy and classical imaging more broadly.
@@ I Introduction
Recently, the field of tensor networks has seen a surge in interest due to their connections to quantum computing. Quantum researchers use tensor networks to design and simulate their quantum machines, while tensor networks experts borrow ideas from quantum theory to emulate and compete with quantum performance Huang et al. (2019); Oh et al. (2023); Liu et al. (2023); Tindall et al. (2024a); Patra et al. (2024). This feedback loop benefits both sides and benefits also scientists and engineers tackling complex real-world problems. To name a few examples, tensor networks have enabled immediate applications in machine learning Stoudenmire and Schwab (2016); Glasser et al. (2018), fluid mechanics Peddinti et al. (2024); Arenstein et al. (2025); van Hülst et al. (2025); Gourianov et al. (2025); Hölscher et al. (2025), finance Rihito et al. (2025), electromagnetism Lively et al. (2025), material science Hauck et al. (2024) and computational chemistry Jolly et al. (2025).
@@
Outline of the paper In the first half of this paper, we focus on natural images and quantum-inspired methods for encoding them into quantum-like objects, which will prove useful for many applications. Several types of matrix product encodings will be compared with one another, along with alternative tensor network structures such as tree tensor networks. We will discuss their compression capabilities and explain their performances in terms of locality, scale separation and entropy using the language of statistical mechanics and quantum physics. The discussion will then shift to classical operations on images represented as tensor network operators, and how these can be applied to classical imaging and image processing. One of the goals of this paper is to draw parallels between Fourier optics and quantum mechanics, and to use these parallels to construct efficient approximations of classical optical propagation in very large dimensions by finding local Hamiltonians whose evolution represents these classical operations. These connections will be often made using ideas stemming from quantum computation—even though no quantum computers are involved—showing how quantum mechanics concepts can lead to efficient classical computations.
II Natural images, resolution and large-scale limit
We define a continuous image (sometimes called a scene or an objet) as the intensity $h$ of an electromagnetic field on a $2d$ plane. From this function of continuous variables, we can consider an approximation on a $2d$ grid of size $L\times L=2^{n}\times 2^{n}$ by evaluating it at discrete points $(x,y)=\Big(\frac{a}{2^{n}},\frac{b}{2^{n}}\Big)$ with $a,b\in\{0,1,2,...,2^{n}-1\}$.
@@ II Natural images, resolution and large-scale limit
As we can not get access to a real continuous image of the physical world (or a model of it), we will have to settle for a digital image as a proxy to the real world and try to separate what comes from the physical scene to what comes from the measurement device (the digital camera here). The question we are interested in here, is how much information and correlation is relevant in an image at the continuous limit and how this can be used to construct an effective model that compress and approximate efficiently the image as well as implement faster optical and imaging simulations on classical computers.
@@
Natural images are closer to real physical systems than they are from random samples of a generic distribution, moreover they are known to be scale invariant with a power spectrum that obeys a power-law distribution Ruderman (1994, 1997). In a more recent study Jobst et al. (2024), the authors showed that, over a large dataset of various natural images of various shapes and forms, the behavior of the Fourier amplitudes follows $\langle\hat{h}(k_{x},k_{y})\rangle\sim\mathcal{O}(k_{x}^{-1.1})\mathcal{O}(k_{y}^{-1.07})$. In this present work, we choose to work with a single “statistically average”image  [^1] As a remark, although the use of “Lena”as the go-to image in the field of image processing has, understandably so, fallen out of favor in recent years for many reasons, its non-power-law behavior could be added to the list of grievances. (cf. Fig. 28) that is very representative of the power-law behavior of natural images. The picture showed numerous times in this paper  [^2] Original image: Roxy.JPEG, dimension $3264\times 2448$, taken with a SONY DSC-F828 camera, sRGB IEC61966-2.1, focal length 38.4 mm, f-number $f/3.2$, exposure time $1/60$ s. follows this behavior honorably well (it is actually closer to $\hat{h}(k)\sim k^{-1}$) and saves us from the time/resource consuming task of averaging over a large dataset of images, with all the generalization and statistical limitations that it entails.
We construct a multi-scale representation of a high-resolution image by iteratively downsampling by a factor of two, starting from a large image of size $L\times L$ down to coarse resolutions Fig. 2. The downsampling is performed using the imresize routine from the ImageTransformations.jl package. Since the procedure involves only coarse-graining operations and no upsampling, it does not introduce spurious degrees of freedom. This argument has been checked numerically, since all tested resizing algorithms have lead to the same results. Each step effectively integrates out short-wavelength fluctuations while preserving long-wavelength structure, in close analogy with a real-space renormalization group (RG) transformation. For natural images, and with standard interpolation, this transformation is well approximated by a low-pass filtering followed by decimation.
From an information-theoretic perspective, the entropy of the image decreases monotonically along this flow, while mutual information between well-separated regions is preserved up to cutoff-scale corrections. Any resampling-induced artifacts are confined to the ultraviolet scale and correspond to RG-irrelevant perturbations that do not affect the large-scale behavior  [^3] Differences between interpolation kernels act as RG-irrelevant operators and are washed out under further coarse-graining. Let us notice that the opposite process of starting with a small image would not satisfied the requirements.. As a result, the sequence of resized images may be regarded as multiple realizations of the same effective system at different linear sizes $L$, allowing one to analyze scale-dependent properties (such as entropy scaling, correlation structure, or compression efficiency) using the standard tools of statistical mechanics in the large-$L$ limit.
Figure 2: Physical interpretation of an image at various resolutions. The downsampling algorithm does not introduce RG relevant features.
The down-sampling procedure used in this work is deterministic, and therefore yields a single representative image at each resolution rather than a full statistical ensemble. As a result, the physical interpretation presented above should be understood in an effective or self-averaging sense, appropriate for large natural images. To make this analogy more explicit, an ensemble at fixed resolution could be generated by introducing randomness only at the ultraviolet scale (for instance by randomizing sub-pixel offsets or local averaging kernels during the coarse-graining step) while preserving long-wavelength structure. Such randomized coarse-graining procedures are analogous to stochastic block-spin renormalization and differ only by RG-irrelevant details. While we do not follow this route in the present work, it represents a necessary path toward a more thorough statistical analysis and will be explored in future research.
III Quantum encodings of images
@@
Next, we are mapping this matrix to a higher dimensional object (a tensor) and then operate compression within this space. The first tensor image encoding that we are considering comes from the field of quantum machine learning, where researchers are interested in efficient ways to feed data to quantum computers. The so-called Flexible Representation of Quantum Images (FRQI) Le et al. (2011) is a quantum encoding scheme that non-trivially represents images by a multi-qubit quantum state $|h\rangle$ by factorizing the pixel values and their positions. More precisely, an image of size $2^{n}\times 2^{n}$ is encoded using a normalized superposition of $2n+1$ qubits
$|\psi\rangle=\frac{1}{2^{n}}\sum_{i=0}^{2^{2n}-1}|c(x_{i})\rangle\otimes|i\rangle,$ (1)
with $|c(x_{i})\rangle=\cos\big(\frac{\pi}{2}x_{i}\big)|0\rangle+\sin\big(\frac{\pi}{2}x_{i}\big)|1\rangle$ [^4] Other kernels are envisageable and well-studied in quantum machine learning, depending on data, context and purposes., where $x_{i}$ encodes the intensity of the pixel at position $i$ (see Fig. 3), and the basis state $|i\rangle$ represents the pixel’s position in the image.
@@
Usually, a row by row flattening of the matrix is used, but more complex paths such as Hilbert, Peano or Moore curves Sagan (2012) can be used. In Fig. 3, for illustration purpose  [^5] It illustrates the fact that any space-filling curve could be in principle be used, a fully-packed self-avoiding walk is generated by a pivot algorithm Oberdorf et al. (2006).
@@
Figure 3: Illustration of matrix to vector reshaping used in FRQI Eq. (1). Below the image, the corresponding vector noted $x_{i}$ with $i=0...2^{2n-1}$.
Intuitively, the path should be as compact as possible such that local correlations in the image remain local upon transformation. In the following, we will use mostly Hilbert curves (see Fig. 4) unless specified otherwise.
Figure 4: Examples of Hilbert curves of order $k$ used for path encoding
@@
This encoding allows for efficient quantum image processing by parallel manipulation of all pixel values through unitary operations on the quantum state. This way, we only need $2n+1$ qubits to store the information of $2^{2n}$ grayscale pixel values, which could lead (in the best cases) to an exponential reduction to the classical case. Although this encoding requires one more qubit that other encodings, it has been shown to work the best on natural images Jobst et al. (2024). The improved representation performance of FRQI over simple amplitude encoding comes from the factorization between spatial degrees of freedom and pixel values inherent to the FRQI construction.
@@
Thanks to the multi-scale nature of typical images, an other interesting and relevant way of encoding images into tensors is the so-called quantics representation Oseledets (2009); Khoromskij (2011) (first appeared in the quantum computing community Wiesner (1996); Zalka (1998); Latorre (2005)) for the tensors as it naturally encodes scale separation. As a result, coarse features are encoded in high-order qubits, fine details live in low-order qubits and scale separation is built into the Hilbert space. It represents a real-space block renormalization expressed in terms of binary degrees of freedom rather than continuous spatial coordinates.
@@
Figure 5: Quantics representation of the vector $\vec{x}$ on $L=8$. Let us denote $\psi_{ijk}$ the resulting tensor components. In this example, we have $x_{1}\rightarrow\psi_{000}$, $x_{2}\rightarrow\psi_{100}$$x_{7}\rightarrow\psi_{111}$ and $x_{8}\rightarrow\psi_{001}$.
@@
Contrary to FRQI, the quantics representation does not flatten the matrix, but factorizes the indices. In $2d$, this means splitting the indices corresponding to spatial coordinates $x$ and $y$ in their binary components $(x_{1}x_{2}x_{3}...x_{n})_{2}$ and $(y_{1}y_{2}y_{3}...y_{n})_{2}$ and rearranging them to form new tensor indices $\mu_{p}=(x_{p}y_{p})_{2}$. This yields a $2n$-order tensor $\psi_{\mu_{1}\mu_{2}...\mu_{n}}$ where each $\mu_{p}$ corresponds to a particular length scale.  [^6] As an example the value of the image at coordinate $(x,y)=(43,78)$ on a grid of size $L=128$ is given by the component $\psi_{00101011001110}$ of a 14-qubits state $|\psi\rangle$. Technically a full tensor of the form $\psi_{\mu_{1}\nu_{1}\mu_{2}\nu_{2}...\mu_{n}\nu_{n}}$ contains $2^{n}\times 2^{n}$ parameters, as many as the initial matrix we are considering, only in a different tensor shape. In both cases (quantics or qubit encoding), we end up with a full tensor that contains exactly the same information as the original image. A non-trivial decomposition of this tensor into an interconnected network might reduce the number of parameters from an exponentially to (in the best cases) a logarithmically small number. This generic approach applied to our original image is illustrated in Fig. 6.
@@
Figure 6: Tensor network compression of a normalized black and white image $h_{ij}$. Here, a generic tensor network is shown for the sake of generality. The approximate symbol will quantify the precision of the compression. The final contracted object is the decompressed image $\tilde{h}_{ij}\approx h_{ij}$.
This tensor network decomposition could be used as a way of compressing images ($h_{ij}\rightarrow\tilde{h}_{ij}$) for transmission and storage or, as we will see later, as a crucial step in a pipeline of transformations allowing to outperform standard optical simulations or image processing. The elephant in the room is that both methods are SVD-based so they require an exponential number of operations to generate the MPS such it might be impractical for very large images $n>30$, the alternative will be to use tensor-cross interpolation. This issue will not be discussed in more detailed but will be one of the focus of attention in the second half of this paper.
@@
Many different topologies of tensor networks Bridgeman and Chubb (2017); Biamonte and Bergholm (2017); Orús (2019) have been proposed over the years and became very powerful theoretical and numerical tools that allow to capture complex phenomena regularly found in the fields of condensed matter, high-energy physics or quantum computation Montangero et al. (2018); Collura et al. (2024). In the following section, we focus on the two most common topologies of tensor networks, namely matrix product states and tensor trees and evaluate their expressive and compressive capabilities for increasing sizes using FRQI and quantics encodings.
IV Matrix Product States and Tensor Tree Networks
@@
The specific tensor network we are focusing on first, is called a matrix product state (MPS) in the physics literature or a tensor train (TT) in the mathematical community.
@@
Figure 7: Illustration of the decomposition of high-order tensor into a matrix product state. The arrow represents a sequential singular value decomposition and $\chi_{k}$ are the bond dimensions.
The full $n$-qubit state vector $|\psi\rangle\in\mathcal{H}$, whether we choose the FRQI or quantics encodings  [^7] Technically the quantics representation does not lead to a proper quantum state while FRQI automatically does, although we can normalize the image before the encoding so we can use the notation $|\psi\rangle$ for both without confusion. is
$|\psi\rangle=\sum_{s_{1},s_{2},...,s_{n}\in\mathbb{Z}_{2}}\psi_{s_{1}s_{2}...s_{n}}|s_{1}s_{2}...s_{n}\rangle,$ (2)
where the tensor $\psi_{s_{1}s_{2}\dots s_{n}}$ is represented by the purple $n$-leg object in Fig. 7. The matrix decomposition outlined below is a generalization of the singular value decomposition and can be used to decompose a generic tensor into constituent tensors of lower order. Every bond corresponds to a Schmidt decomposition of the Hilbert space $\mathcal{H}=\mathcal{H}_{[1,2,...,k]}\otimes\mathcal{H}_{[k+1,k+2,...,n]}$. This succession ($\it{cf.}$ Appendix A for details) of operations leads to the usual form of the factorization of components
$\psi_{s_{1}s_{2}\dots s_{n}}=\sum_{\alpha_{1},\dots,\alpha_{n-1}}A^{[1]}_{s_{1},\alpha_{1}}A^{[2]}_{\alpha_{1},s_{2},\alpha_{2}}..A^{[n]}_{\alpha_{n-1},s_{n}}$, (3)
which is represented diagrammatically by the black object in Fig. 7. Depending on the scaling of $\chi_{k}$ with $n$, the MPS can efficiently represent $\psi_{s_{1}s_{2}\dots s_{n}}$. This point will be the focus on the next section.
@@
Once the image is encoded into the network, the numerical evaluation of the compressive power versus the reconstruction error is measured and the total number of parameters in the network is called the total dimension $D$. Now, we apply the decomposition to the picture showed earlier and analyze the error of the decomposition with respect to the original image for increasing allowed rank and increasing size. The goal is to mirror what we would have done with a typical stat-mech system, by studying the rank of the MPS for various sizes and infer the large-scale limit.
Instead of using the quantum fidelity to evaluate the quality of the reconstruction, we define the mean squared error (as it is more common in machine learning) between the original matrix $h_{ij}$ and the decompressed image $\tilde{h}_{ij}$ as
@@
Figure 8: Scaling of the mean squared error of the compression depending on system size for Hilbert-FRQI MPS and quantics tensor trains ($q\mathrm{TT}$).
Additionally, we introduce the inverse compression ratio as $\xi(L)=DL^{-2}$, which is the relevant order parameter that quantifies the efficiency of the approximation. Numerical simulations (see Fig. 8) on qubit encoding matrix product states and quantics tensor trains suggest the following finite-size scaling
@@
with $\phi(\xi)\sim\xi^{-\alpha}$ with $\alpha\approx 0.75$ and $\kappa\approx 1.5$ for $\xi\ll 1$. In this efficient regime $\epsilon(L)\sim\mathcal{O}(1)$, meaning that the error is (modulo multiplicative logarithmic corrections) independent of system size (consistent with Jobst et al. (2024)), and results are fairly similar between the two encodings. Conversely, the compression ratio at fixed error increases polynomially $\epsilon=C^{0.75}\times L^{1.5}$, so $C(L)\sim L^{2}$. This is in sharp contrast with local compressors like JPEG, since they possess a maximum compression ratio which is given by the division of the original image in small finite-size squares, this point will be made clearer later on.
@@
We define $\xi(L)=1$ as the compressive threshold, as for $\xi(L)<1$ the approximation is compressive, while for $\xi(L)>1$ the number of parameters becomes greater than the original image. In the overfitting regime (the red region $\xi(L)\geq 1$), the error goes to zero exponentially quickly at a value $\xi\rightarrow\xi_{\infty}$. We notice that $\xi_{\infty}^{q\mathrm{TT}}<\xi_{\infty}^{\mathrm{MPS}}$ by a big margin (because there is one less qubit needed in the quantics representation than FRQI). This is due to the fact that, by construction of the singular value decomposition, the dimension of a random MPS of $n$ sites can be showed recursively to be $D_{\mathrm{max}}(n)=8/3\times 2^{n}$. This can be found by integrating the exponential Page’s curve Page (1993) of a Haar state over every bonds Fig. 9. It implies $\xi_{\infty}^{q\mathrm{TT}}=8/3$ and $\xi_{\infty}^{\mathrm{MPS}}=16/3$ and at very high resolution the error is mostly dominated by pixel noise in the image.
Figure 9: Evolution of the bond dimensions along the $q\mathrm{TT}$ for $n=20$ for $\epsilon=10^{-1},10^{-2},...,10^{-12}$ (bottom to top curves). Scale separation is very noticeable at large error, while the ranks are maximized (Haar-like) for lossless reconstruction.
The linear network treats noise as global, since every bond is saturated. Although interleaved quantics and FRQI encodings based on Hilbert curves differ at the level of microscopic ordering, both induce a hierarchical dyadic decomposition of the image and preserve locality across scales. As a result, they lead to equivalent entanglement scaling and compression behavior when used with linear tensor networks. Interestingly, the value of $\xi_{\infty}$ is also of certain relevance in the intermediate regime (black region in Fig. 8) near the compressive threshold $\xi=1$ since a smaller value $\xi_{\infty}$ seems to force the curve to accelerate faster towards this line [^8] A more fondamental question is whether or not a real image of the physical world (not a digital picture) would admit a true area-law at any reasonable scale.. This fact allows the $q\mathrm{TT}$ to improve over the Hilbert MPS in this intermediate region. This argument will be made more clear in the following section.
@@
Another important decomposition is the binary tree tensor network (TTN) where the constituent tensors of the network lie on the nodes of a connected acyclic tree graph. The appeal of tensor tree network comes down to their relative ease of constructions and contractions as well as their greater expressive power and has been recently explored in the context of multivariate function approximations Tindall et al. (2024b).
@@
Figure 10: Left: Binary tensor tree network $\mathrm{TTN}_{B}$ with 14 physical bulk sites.
The construction of tree tensor networks is not excessively different than the case of matrix product states. Once fixed a binary tree topology, every SVD corresponds to a tree cut which can separate non-contiguous qubits in the original ordering $\{1,2,...,k\}|\{k+1,k+2,...,n\}$. We can decompose the full tensor into a binary tree of lower-order tensors using a sequence of Schmidt decompositions corresponding to the cuts defined by the edges of the tree. Every edge $e$ of the tree defines a bipartition of the Hilbert space
$\mathcal{H}=\mathcal{H}_{\text{subtree}(e)}\otimes\mathcal{H}_{\text{rest}(e)}.$ (6)
The construction is detailed in appendix A and a factorization of each subtree recursively into tensors associated with each node of the binary tree produces the usual tree decomposition of the tensor components
$\psi_{s_{1}s_{2}\dots s_{n}}=\sum_{\{\alpha_{e}\}}\prod_{v\in\text{nodes}}A^{(v)}_{s_{v},\alpha_{v_{1}},\alpha_{v_{2}},\alpha_{v_{\text{parent}}}}$, (7)
where each $A^{(v)}$ has indices for the physical qubit at node $v$, the virtual bonds to its children, and the bond to its parent (if any). This is the direct generalization of the MPS canonical form to a binary tree geometry.
@@
For notation convenience, we introduce the notation $q\mathrm{TTN}_{B}$ for quantics encoding on this tree. In Fig. 11, we plot the scaling of $\xi(L)$ for $q\mathrm{TTN}_{B}$ and observe a similar scaling as before. In quantum many-body physics, it is common knowledge that matrix product states can efficiently represent $1d$ quantum systems with area-law entanglement. In contrast, the hierarchical structure of tensor trees can also describe more complex types of entanglement, including some long-range quantum correlations, in higher-dimensional or critical systems where entanglement scaling can exceed that of an area law. In the scenario that we are interested in, the image considered follows an area law (since the bond dimensions of the MPS remain constant up to quite a large precision), and tensor trees only marginally improve the compression ratio but not the scaling of the function $\xi(L)$.